# NEST: Self-supervised Fast Conformer as All-purpose Seasoning to Speech Processing Tasks

### v2

@@ Abstract
Self-supervised learning has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of these approaches are computationally intensive due to using transformer encoder and lack of sub-sampling. In this paper, we propose a new self-supervised learning model termed as pretraiNed Encoder for Speech Tasks (NEST). Specifically, we adopt the FastConformer architecture, which has an 8x sub-sampling rate and is faster than Transformer or Conformer architectures. Instead of clustering-based token generation, we resort to fixed random projection for its simplicity and effectiveness. We also propose a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that the proposed NEST improves over existing self-supervised models on a variety of speech processing tasks. Code and checkpoints will be publicly available via NVIDIA NeMo toolkit [^1] https://github.com/NVIDIA/NeMo.
@@ I Introduction
Figure 1: NEST serves as a bird nest that incubates the variety of speech task models.
@@
Figure 2: (a) The proposed NEST architecture. (b) Two ways to use NEST encoder: (left) use as weight initialization for tasks that require more parameters (e.g., speech recognition); (right) learn weighted summation of features from different layers of the frozen NEST for tasks that require less trainable parameters (e.g., speaker verification).
@@
In this paper, we tackle all these challenges and bring the best practices from previous works, which constitute the proposed pretraiNed Encoder for Speech Tasks (NEST) framework. Our contributions are summarized as follows:
@@ II Approach
In this section, we introduce the proposed NEST framework for learning efficient self-supervised speech encoder, as illustrated in Figure 2(a).
@@ II-C Speech Quantization
We resort to BEST-RQ [7] as our speech quantization method. Specifically, we use a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectrogram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens.
@@ II-E Training
Since quantization is performed before the convolutional sub-sampling, there is a mismatch in the lengths between the predicted tokens and target tokens. To mathc the sequence lengths, target tokens are averaged for every 8 frames, then apply threshold of 0.9 to select frames to be taken into loss calculation. Cross-entropy loss is applied on selected tokens determined by the input masks.
@@ III-A Dataset and Settings
We train the NEST-L (108M) and NEST-XL (600M) models using 100K hours of both English speech data, including 60K hours from LibriLight [22], 24K hours from English subset of Voxpopuli [23], and about 20K hours sampled data from the combination of Fisher [24], Switchboard [25], WSJ [26], NSC [27], People’s Speech [28]. The audios for speech augmentation is randomly selected within each batch, while we use noise audios from MUSAN [29] and Freesound [30]. We train the models with global batch size of 2048 for 80K steps on 128 NVIDIA A100 GPUs, with Noam annealing [16] and peak learning rate of 0.004, weight decay of 1e-3, gradient clipping 1.0 and warm-up of 25K steps. We set the speech augmentation probability as 0.2, among which we set noise and speech augmentation probabilities as 0.1 and 0.9 respectively. Code and checkpoint will be publicly available.
@@ III-B Results on SUPERB Multi-task Speech Processing
We evaluate our model’s performance on the SUPERB [18] benchmark for multi-task evaluation on self-supervised speech models. For speech recognition (ASR), phoneme recognition (PR) and speaker diarization (SD) tasks, we use the architecture in the left part of Figure 2(b) and a simple linear layer as the task decoder. We train ASR and PR with CTC [31] loss, while the SD task is trained with permutation invariant loss [32]. For speaker identification/verification (SID/SV), keyword spotting (KS) and emotion recognition (ER) tasks, we resort to the architecture presented in the right part of Figure 2(b), and use the ECAPA-TDNN-small [33] as the task decoder. We following the same train/val/test splits as in the SUPERB [18] and train the models for 100 epochs.
@@
TABLE IV: DER results on speaker diarization. EEND-GLA [34] applies additional clustering after end-to-end diarizer, so we exclude it from the comparison for fairness. “*” indicates second best.
@@
EEND-GLA-small [34] 14.39 6.94 11.42 14.49
EEND-EDA [35] 15.55 7.83 12.29 17.59
NeMo MSDD [36] 29.40 11.41 16.45 19.49
RandFC-L-Linear 22.68 12.17 19.16 24.47
NEST-L-Linear 19.63 8.56 15.85 24.73
NEST-L-Sortformer [sortformer] 17.06* 6.91* 10.34* 23.07
@@ III-C Results on Multi-lingual Speech Recognition
Besides multi-task evaluation, we also study if an SSL model trained on single language can help other languages. To this end, we train an ASR model on four different languages: English (En), German (De), French (Fr), Spanish (Es). Specifically, we train an ASR model using NEST-XL as weight initialization and the hybrid-CTC-RNNT loss [37]. The training data comprises of 2.5K hours of German speech (MCV, MLS, Voxpopuli), 8.5K hours of English speech (MCV [38], MLS [39], Voxpopuli [23], SPGI [40], Europarl [41], LibriSpeech [42], NSC1 [27]), 1.4K hours of Spanish speech (MCV, MLS, Voxpopuli, Fisher [24]) and 1.9K hours of French speech (MCV, MLS, Voxpopuli). For baselines, we train another model using an English ASR model [43] as weight initialization, and also include some of the best ASR models like Whisper [20], SeamlessM4T [19] and Canary [21]. We run all models with the same beam size 5 without extra language models on test sets of MCV-16.1 [38] and Voxpopuli [23].
@@ III-E Results on Speaker Diarization
To evaluate our model’s performance on speaker diarization, we follow the Sortformer [sortformer] architecture and attach 18 layers of transformer encoder [16] after the NEST encoder. We also train a simpler baseline with only FastConformer encoder and a linear decoder, with either random initialization or NEST initialization. For training data, we use a combination of real data (Fisher [24]) and simulated data (composed from LibriSeech [42], SRE [47, 48]) generated by the NeMo speech data simulator [49], with a total of 6.4K hours. We compare with state-of-the-art diarization models EEND-EDA [35] and EEND-GLA [34] and show the results in Table IV.
By comparing RandFC-L-Linear and NEST-L-Linear, we can see that NEST is able to significantly reduce the DER of a simple FC-Linear baseline, showing the importance of using NEST in speaker diarization. We can also notice that Sortformer [sortformer] with NEST initialization is able to achieve second best results on three of the four evaluated test sets, which are comparable to the state-of-the-art EEND-EDA [35].
@@ IV Conclusion
In this paper, we propose a self-supervised model NEST , which is faster than previous SSL models by applying FastConformer [17] as the speech encoder. We simplify the audio quantization by applying fixed random-projection quantization [7], and also generalize the noisy speech augmentation [8] to allow for more randomness and speakers. Extensive experiments on multiple speech processing tasks show that the NEST model can help achieve state-of-the-art performance. Code, configurations and checkpoints are also publicly available through NVIDIA NeMo toolkit [56].

### v3

@@ Abstract
Self-supervised learning has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of current approaches are computationally expensive due to lacking in sub-sampling or using clustering based speech quantization. In this paper, we propose a simplified and more efficient self-supervised learning framework termed as NeMo Encoder for Speech Tasks (NEST). Specifically, we adopt the FastConformer architecture, which has an 8x sub-sampling rate and is faster than Transformer or Conformer architectures. Instead of clustering-based token generation, we resort to fixed random projection for its simplicity and effectiveness. We also propose a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that NEST improves over existing self-supervised models on a variety of speech processing tasks. Code and checkpoints will be available via NVIDIA NeMo toolkit [^1] https://github.com/NVIDIA/NeMo.
@@ I Introduction
Figure 1: (a) The proposed NEST framework for speech self-supervised learning. (b) Two ways to use NEST encoder: (left) use as weight initialization for tasks that require more parameters (e.g., speech recognition); (right) learn weighted summation of features from different layers of the frozen NEST for tasks that require less trainable parameters (e.g., speaker verification).
@@
In this paper, we tackle all these challenges and bring the best practices from previous works, which constitute the proposed NeMo Encoder for Speech Tasks (NEST) framework. Our contributions are summarized as follows:
@@ II Approach
The proposed NEST framework is illustrated in Figure 1(a).
@@ II-C Speech Quantization
We resort to BEST-RQ [7] as our speech quantization method. Specifically, we use a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectrogram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens. Since there is a 8x subsampling, we channel-concatenate the features for each 8 consecutive frames before feeding into the linear layer, such that the lengths for the target tokens and input features are equal.
@@ II-E Training
Since masking is performed before the convolutional sub-sampling, there is a mismatch in the lengths between the predictions and masks. To match the sequence lengths, masks are averaged for every 8 frames, then apply threshold of 0.9 to select frames to be taken into loss calculation. Cross-entropy loss is applied on selected positions determined by the averaged input masks.
@@ III-A Dataset and Settings
We train the NEST-L (108M) and NEST-XL (600M) models using 100K hours of both English speech data, including 60K hours from LibriLight [22], 24K hours from English subset of Voxpopuli [23], and about 20K hours sampled data from the combination of Fisher [24], Switchboard [25], WSJ [26], NSC [27], People’s Speech [28]. The audios for speech augmentation is randomly selected within each batch, while we use noise audios from MUSAN [29] and Freesound [30]. We train the models with global batch size of 2048 for 80K steps on 128 NVIDIA A100 GPUs, with Noam annealing [16] and peak learning rate of 0.004, weight decay of 1e-3, gradient clipping 1.0 and warm-up of 25K steps. We set the speech augmentation probability as 0.2, among which we set noise and speech augmentation probabilities as 0.1 and 0.9.
@@ III-B Results on SUPERB Multi-task Speech Processing
We evaluate our model’s performance on the SUPERB [18] benchmark for multi-task evaluation on self-supervised speech models. For speech recognition (ASR), phoneme recognition (PR) and speaker diarization (SD) tasks, we use the architecture in the left part of Figure 1(b) and a simple linear layer as the task decoder. We train ASR and PR with CTC [31] loss, while the SD task is trained with permutation invariant loss (PIL) [32]. For speaker identification/verification (SID/SV), keyword spotting (KS) and emotion recognition (ER) tasks, we resort to the architecture presented in the right part of Figure 1(b), and use the ECAPA-TDNN-small [33] as the task decoder. We following the same train/val/test splits as in the SUPERB [18] and train the models for 100 epochs.
@@
Model Params Data (hrs) Europarl mExpresso Fleurs Average
@@
TABLE IV: DER results on speaker diarization. “*” indicates second best.
@@
EEND-EDA [34, 35] 15.55 7.83* 12.29 17.59
NeMo MSDD [36] 29.40 11.41 16.45 19.49*
RandFC-L-MLP 21.71 11.60 15.89 24.34
NEST-L-MLP 16.83 7.88 11.71* 23.46
RandFC-L-Sortformer 18.93 9.39 13.56 23.30
NEST-L-Sortformer [sortformer] 16.06* 6.40 11.02 20.10
@@ III-C Results on Multi-lingual Speech Recognition
Besides multi-task evaluation, we also study if an SSL model trained on single language can help other languages. To this end, we train an ASR model on four different languages: English (En), German (De), French (Fr), Spanish (Es). Specifically, we train an ASR model using NEST-XL as weight initialization and the hybrid-CTC-RNNT loss [37]. The training data comprises of 2.5K hours of German speech (MCV, MLS, Voxpopuli), 8.5K hours of English speech (MCV [38], MLS [39], Voxpopuli [23], SPGI [40], Europarl [41], LibriSpeech [42], NSC1 [27]), 1.4K hours of Spanish speech (MCV, MLS, Voxpopuli, Fisher [24]) and 1.9K hours of French speech (MCV, MLS, Voxpopuli). For baselines, we train another model using an English ASR model [43] as weight initialization, and also include some of the best ASR models like Whisper [20], SeamlessM4T [19] and Canary [21]. We run all models with the same beam size 5 with no language models on test sets of MCV-16.1 [38] and Voxpopuli [23].
@@ III-E Results on Speaker Diarization
To evaluate our model’s performance on speaker diarization, we train two variants of end-to-end diarization models: (1) simple two-layer multi-layer-perceptron (MLP) on top of FastConformer encoder and train with PIL; (2) more sophisticated Sortformer [sortformer] model with 18 layers of transformer on top of the encoder and train with a hybrid PIL and sorting loss. We also apply both NEST and random initialization to both methods for comparison. For training data, we use a combination of real data (Fisher [24], AMI Mix-Headset [47], ICSI [48], DIHARD3 dev [49], VoxConverse v0.3 [50], AISHELL [51]) and simulated data (composed from LibriSpeech [42], SRE [52, 53]) generated by the NeMo speech data simulator [54], with a total of 6.4K hours. We evaluate models’ performance on DIHARD3-eval [49] and CALLHOME-part2222We used splits from the Kaldi x-vector recipe [55]: Part 1 for training and Part 2 for evaluation [56].
As in Table IV, by comparing RandFC-L-MLP with NEST-L-MLP, and RandFC-L-Sortformer with NEST-L-Sortformer, we can see that NEST is able to significantly reduce the DER of models using randomly initialized encoder, showing the importance of using NEST in speaker diarization. We can also notice that Sortformer [sortformer] with NEST initialization is able to achieve second best results on DIHARD3 eval set, and outperforms EEND-EDA [34] on 2 and 3 speaker settings of the CALLHOME-part2.
@@ IV Conclusion
In this paper, we propose a speech self-supervised framework termed NEST, and extensive experiments on multiple speech processing tasks show that the NEST model can help achieve state-of-the-art performance. Code, configurations and checkpoints are available through NVIDIA NeMo toolkit.